OpenStudy (anonymous):

Please help!!!! A sample containing 2.50 moles of an ideal gas at STP (0 degrees celsius, 760 torr) would have a volume of what in __________ Liters?

OpenStudy (anonymous):

Looks like a plug-and-play problem. You just need the equation. Ideal gas law problem?

OpenStudy (anonymous):

pV=nRT

OpenStudy (anonymous):

@Kinzan pressure 760 torr x volume ? = n x r x temperature 273.15 kelvin what does N and R stand for?

OpenStudy (abmon98):

760 torr when changed to pascals is equal to 101325(atmospheric pressure) so apply in the equation that says pV=nRT n stands for the number of mole in which youve got 2.5 and R is a gas constant value which is 8.3 so 101325*v=2.5*8.3*273

OpenStudy (anonymous):

@Abmon98 so by multiplying 2.5*8.3 * 273 = 5664.75

OpenStudy (abmon98):

divide this number by 101325 and multiply it by 1000 because you want the units in liters

OpenStudy (anonymous):

@Abmon98 see i first did this 760 torr * ? =2.50 moles * 0.08205 * 273 = 56.0

OpenStudy (abmon98):

gas constant is not equal to 0.08205